OpenStudy (anonymous):

Answer For Medal Show how changing the position of the parentheses affects the solution for the expression: forty divided by 5 plus 5 times the quantity of 2 to the second power minus 1. Show all of your work

OpenStudy (anonymous):

Please answer this I only have 5 mins left

OpenStudy (anonymous):

Hi, Let me try 1.) 54.5 -08.7 ------------------ 45.8 2.) 9+6 = 15 15 X 6 --------- = 11.25 => 11.25^3 = 1423.8 8 3.) 56.70 + 56.08 ------------- 0.1 the answer is eaqual to 0.1 4.) 46.9 ----------- = 469 10 469 5.3 --------- X ------- = ------ = 8.84 10 5.3 53 5.) 7.5 + 2.5 + 4.9 = 14.9

OpenStudy (anonymous):

is that how you explain it because I need this to show the work

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome